OpenStudy (anonymous):

Which of the following is the solution to the equation 8n = 125(n + 4) ? Answer n = 12 n = –4 n = 4 n = –12

OpenStudy (anonymous):

is the problem written correctly?

OpenStudy (anonymous):

ahhh no its actually sqaureroot of 5^8n=125^(n+4)

OpenStudy (pfenn1):

Is this what you mean?\[\sqrt{5^{8n}}=125^{(n+4)}\]